# A Multi-way Parallel Named Entity Annotated Corpus for English, Tamil and Sinhala

### v1

@@ 7 Case Study - Improving Neural Machine Translation with NER
Named Entity translation is challenging even to the modern day NMT systems [Läubli et al., 2020]. Early solutions to this problem involve the use of external transliteration systems [Grundkiewicz & Heafield, 2018, Ameur et al., 2017, Zhang et al., 2020]. However, such NE transliteration systems focus on translating individual NEs with no regard to the context of the NEs, which makes it difficult to resolve ambiguities. This is particularly problematic in highly inflected languages such as Sinhala and Tamil. Hu et al. [2022] recently proposed an alternative solution that involves pre-training a language model prior to fine-tuning it for NMT. Since their results seem to far exceed the results of NMT models that do not have explicit NE translation capabilities, we selected this model as the case study to demonstrate the usability of our NER system.
@@
Hu et al. [2022] adopted a pre-training and fine-tuning process to implement an NMT system that has better capabilities for NE translation. They first identified NEs in a monolingual corpus and linked them to a knowledge base (KB) that contains entity translations (WikiData [Vrandečić & Krötzsch, 2014] in their case). For NE linking, they used the SLING [Ringgaard et al., 2017] entity linker. Then the entity translations in the KB were used to generate noisy code-switched data. After that, a Transformer model [Vaswani et al., 2017] was pre-trained with this noisy code-switched data using the de-noising pre-training objective. Note that this is the same objective that was used to train the encoder-decoder pLM mBART [Tang et al., 2021]. In de-noising pre-training, the Transformer model is taught to reconstruct an original sentence from its noised version. Finally to further improve translation of low-frequency NEs, they used a multi-task learning strategy that fine-tunes the model using both the denoising task on the monolingual data and the translation task on the parallel data. This model, known as DEEP (DEnoising Entity Pre-training) is illustrated in Fig. 3 using English and Sinhala as the example (note the code-switched data used for pre-training).
Figure 3: DEEP[Hu et al., 2022] Architecture
@@
We first implemented Hu et al. [2022]’s DEEP NMT system using SLING, as well as a baseline NMT system without DEEP, for English-Sinhala (En-Si) translation. We used WikiData [Vrandečić & Krötzsch, 2014] as the monolingual dataset, and Fernando et al. [2020]’s parallel data to train the NMT model. To elaborate further, the baseline NMT system was built by simply fine-tuning a Transformer model from scratch, similar to Hu et al. [2022]. To build the DEEP NMT system, first the NEs in the Sinhala WikiData were identified and linked to their corresponding English NEs using SLING. Then the NEs in the Sinhala WikiData were replaced with the identified English NEs. These code-switched noisy sentences were used to pre-train a Transformer model using the de-noising objective. Finally, this pre-trained Transformer model was further fine-tuned with the training set of  Fernando et al. [2020]’s parallel data using the NMT objective.
@@
Next, we re-implemented the DEEP system by replacing SLING with our best NER system. In other words, we used our NER system to identify NEs in Sinhala WikiData. Entity linking across the languages was done using the Pywikibot library [pfu, 2015]. As shown in Table 7, the resulting NMT system significantly outperforms both baseline and DEEP with SLING (by about 9 BLEU points).
@@ 8 Conclusion
This paper presented a multi-way parallel Named Entity annotated dataset for Sinhala-English-Tamil. According to the best of our knowledge, this is the first multi-way parallel NE annotated corpus for any language pair. We carried out a comprehensive evaluation of different types of pre-trained pLM for the task of Named Entity Recognition. We present our multilingual NER model, which was trained with the multi-way parallel corpus as our best pick for the task. We also demonstrated the utility of our NER system, by plugging it into an English-Sinhala NMT model.

### v2

@@ 7 Case Study - Improving Neural Machine Translation with NER
Named Entity translation is challenging even to the modern day NMT systems [64]. Early solutions to this problem involve the use of external transliteration systems [65, 66, 67]. However, such NE transliteration systems focus on translating individual NEs with no regard to the context of the NEs, which makes it difficult to resolve ambiguities. This is particularly problematic in highly inflected languages such as Sinhala and Tamil. Hu et al. [2] recently proposed an alternative solution that involves pre-training a language model prior to fine-tuning it for NMT. Since their results seem to far exceed the results of NMT models that do not have explicit NE translation capabilities, we selected this model as the case study to demonstrate the usability of our NER system.
@@
Hu et al. [2] adopted a pre-training and fine-tuning process to implement an NMT system that has better capabilities for NE translation. They first identified NEs in a monolingual corpus and linked them to a knowledge base (KB) that contains entity translations (WikiData [68] in their case). For NE linking, they used the SLING [69] entity linker. Then the entity translations in the KB were used to generate noisy code-switched data. After that, a Transformer model [70] was pre-trained with this noisy code-switched data using the de-noising pre-training objective. Note that this is the same objective that was used to train the encoder-decoder pLM mBART [71]. In de-noising pre-training, the Transformer model is taught to reconstruct an original sentence from its noised version. Finally to further improve translation of low-frequency NEs, they used a multi-task learning strategy that fine-tunes the model using both the denoising task on the monolingual data and the translation task on the parallel data. This model, known as DEEP (DEnoising Entity Pre-training) is illustrated in Fig. 3 using English and Sinhala as the example (note the code-switched data used for pre-training).
Figure 3: DEEP[2] Architecture
@@
We first implemented Hu et al.’s [2] DEEP NMT system using SLING, as well as a baseline NMT system without DEEP, for English-Sinhala (En-Si) translation. We used WikiData [68] as the monolingual dataset, and Fernando et al.’s [57] parallel data to train the NMT model. To elaborate further, the baseline NMT system was built by simply fine-tuning a Transformer model from scratch, similar to Hu et al. [2]. To build the DEEP NMT system, first the NEs in the Sinhala WikiData were identified and linked to their corresponding English NEs using SLING. Then the NEs in the Sinhala WikiData were replaced with the identified English NEs. These code-switched noisy sentences were used to pre-train a Transformer model using the de-noising objective. Finally, this pre-trained Transformer model was further fine-tuned with the training set of Fernando et al.’s [57] parallel data using the NMT objective.
@@
Next, we re-implemented the DEEP system by replacing SLING with our best NER system. In other words, we used our NER system to identify NEs in Sinhala WikiData. Entity linking across the languages was done using the Pywikibot library [72]. As shown in Table 7, the resulting NMT system significantly outperforms both baseline and DEEP with SLING (by about 9 BLEU points).
@@ 8 Conclusion
This paper presented a multi-way parallel Named Entity annotated dataset for Sinhala-English-Tamil. According to the best of our knowledge, this is the first multi-way parallel NE annotated corpus for any language pair. We carried out a comprehensive evaluation of different types of pre-trained pLM for the task of Named Entity Recognition. We present our multilingual NER model, which was trained with the multi-way parallel corpus as our best pick for the task. We also demonstrated the utility of our NER system, by plugging it into an English-Sinhala NMT model. In future, we plan to increase the size of the dataset further and extend the annotation to cover more NEs.
@@ Acknowledgements
Declaration
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. Generative AI tools were not used to prepare the manuscript. The authors do not have anything else to declare.